OpenStudy (anonymous):

What is the smallest positive value for x where y=sin2x reaches its maximum? How do you figure this out?

OpenStudy (anonymous):

solve \[\sin(2x)=1\]since that is the biggest sine can be

OpenStudy (anonymous):

you should get \[2x=\frac{\pi}{2}\implies x=\frac{\pi}{4}\]